OpenStudy (anonymous):

Multiply and simplify: -2x^2 * x^3y A. -2x^5y^3 B. -2x^6y C. -2x^5y D. 4x^5y

OpenStudy (anonymous):

any ideas for this one? when you multiply, if the bases are the same, you add the exponents

OpenStudy (anonymous):

...

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

Are you sure that questions and choices are true

OpenStudy (anonymous):

no, not D

OpenStudy (anonymous):

\[-2x^2 \times x^3y=-2\times x^{2+3}y=-2x^5y\] \]

OpenStudy (anonymous):

so its C then